The production function is Y= 2K^0.5•N^0.5. Where the real wage (w) is 4,
what is the capital per worker (k) in the economy?
Answer #1

Marginal product of labor = Wage

Marginal product of labor , MPN = (2)(0.5) K0.5 N-0.5

= K0.5 N-0.5

= K0.5 / N0.5 = k0.5 (because K/N =k represents capital per worker)

At optimal MPN = w

k0.5 = 4 (because w=4 (given) )

k = (4)2

k = 16

Hence, capital per worker (k) in the economy is 16 units.